OpenStudy (anonymous):

x+2 * 4x ---- ---- multiply -2 q +5

OpenStudy (lgbasallote):

multiply the numerator; multiply the denominator

OpenStudy (lgbasallote):

\[\huge \frac{x+2}{-2} \times \frac{4x}{q+5} \implies \frac{ 4x(x+2)}{-2(q+5)}\] simplify to get the final answer

OpenStudy (lgbasallote):

dos that help?

OpenStudy (anonymous):

maybeh :)

OpenStudy (anonymous):

thank you

OpenStudy (lgbasallote):

welcome ^_^